# Neural Context Flows for Meta-Learning of Dynamical Systems

### v5

@@ 1 Introduction
    We present an interpretable framework for Multi-Task representation learning, incorporating a straightforward method for uncertainty estimation. This work extends the emerging trend of explainable linearly parameterized physical systems Blanke & Lelarge (2024) to non-linear settings, thus broadening its applicability. For affine systems, we provide a lean proof for the identifiability of their underlying parameters.
@@ 2 Related Work
While LEADS excels at interpolation tasks, it performs poorly during extrapolation Kirchmeyer et al. (2022). Furthermore, it requires retraining a new network $g^{e}_{\psi}$ each time a new environment is encountered, which can be constraining in scenarios where adaptation is frequently required. Before LEADS, other MTL approaches had been proposed outside the context of dynamical systems Caruana (1997); Rebuffi et al. (2017; 2018); Lee & Parish (2021). Still, they do not address the crucial adaptation to new tasks which is our focus.
@@ 3 Neural Context Flow
We depict the NCF framework in Fig. 2, along with a 3-networks architecture that lifts the contexts and the state vectors into the same representational space before they can interact. Concretely, $x_{i}^{e}(t)$ and the context $\xi$ are first processed independently into $\tilde{x}_{i}^{e}(t)$ (by a state network in blue) and $\tilde{\xi}$ (by a context network in green) respectively, before they are concatenated and fed to a main network (in purple) to produce $f_{\theta}(x_{i}^{e}(t),\xi)$. This explicitely allows the model to account for the potentially nonlinear relationship between the context and the state vector at each evaluation of the vector field.
@@ 3.1 Meta-Training
The above comparison demands the definition of two variants of Neural Context Flows. NCF-$\mathit{t}_{2}$ uses second-order Taylor expansion in equation 5, and is trained via proximal alternating minimization (Algorithm 1). NCF-$\mathit{t}_{1}$ on the other hand, is implemented using a first-order Taylor expansion, and trained using ordinary alternating minimization (Algorithm 3). Although less expressive than NCF-$\mathit{t}_{2}$, NCF-$\mathit{t}_{1}$ is faster and serves as a powerful baseline in our experiments.
@@
2:$\,\,\theta_{0}\in\mathbb{R}^{d_{\theta}}$ randomly initialized
3:$\,\,\xi^{1:m}_{0}=\bigcup\limits_{e=1}^{m}\xi^{e}$, where $\xi^{e}=\mathbf{0}\in\mathbb{R}^{d_{\xi}}$
4:$\,\,q_{\text{max}}\in\mathbb{N}^{*}$; $\beta\geq L\in\mathbb{R}^{+};\eta_{\theta},\eta_{\xi}>0$
@@ 4.1 Experimental Setting
The NCF framework is evaluated on five seminal benchmarks. The Simple Pendulum (SP) models the periodic motion of a point mass suspended from a fixed point. The Lotka-Volterra (LV) system models the dynamics of an ecosystem in which two species interact. Additional ODEs include a simple model for yeast glycolysis: Glycolytic-Oscillator (GO) Daniels & Nemenman (2015), and the more advanced Sel’kov Model (SM) Strogatz (2018); Sel’Kov (1968). Like GO, SM is non-linearly parameterized, but additionally exhibits starkly different behaviors when its key parameter is varied (see LABEL:fig:attractors). Finally, we consider three PDEs, all with periodic boundary conditions, and cast as ODEs via the method of lines: the non-linear oscillatory Brusselator (BT) model for autocatalytic chemical reactions Prigogine & Lefever (1968), the Gray-Scott (GS) system also for reaction-diffusion in chemical settings Pearson (1993), and Navier-Stokes (NS) for incompressible fluid flow Stokes et al. (1851).
@@ 4.2 Interpolation Results
Train (×$\times 10^{-1}$) Adapt (×$\times 10^{-3}$)
@@
Table 1 emphasize the adaptation-time merits of Meta-Learning approaches like NCFs in lieu of baselines where one context-agnostic vector field is trained for all environments indiscriminately (One-For-All or OFA), or for each environment independently (One-Per-Env or OPE). Additional results and further analysis of these differences is done in Section B.3.
@@ 4.3 Extrapolation Results
A Meta-Learning algorithm is only as good as its ability to extrapolated to unseen environments. In this section we tackle several one-shot generalization problems, i.e. $S^{\prime}=1$. We consider the LV, GO, SM, GS, BT, and NS problems, which all involve adaptation environments outside their meta-training distributions. We compare both variants of NCF to two baselines: CAVIA Zintgraf et al. (2019) which is conceptually the closest GBML method to ours, and CoDA-$\ell_{1}$ Kirchmeyer et al. (2022). Their hyperparameters, laid out in Appendix D, are tuned for a balance of computational efficacy and performance, all the while respecting each baseline’s key assumptions.
At similar parameter counts [^3] This accounts for the total number of learnable parameters in the framework excluding the context vectors, whose size might vary with the method as explained in Appendix D., Table 2 shows that CAVIA is the least effective for learning all six physical systems, with a tendency to overfit on the few-shots it receives during both meta-training and meta-testing Mishra et al. (2017). NCF-$\mathit{t}_{2}$ achieves SoTA OoD results on 5 out of 6 problems. While CoDA retains its superiority on GS, we find that it struggles on non-linear problems, most notably on the SM problem whose trajectories go through 3 distinct attractors in a Hopf bifurcation.
@@
CAVIA 305246 91.0$\pm$63.6 120.1$\pm$28.3 130711 64.0$\pm$14.1 463.4$\pm$84.9
CoDA 305793 1.40$\pm$0.13 2.19$\pm$0.78 135390 5.06$\pm$0.81 4.22$\pm$4.21
NCF-$\mathit{t}_{1}$ 308240 6.73$\pm$0.87 7.92$\pm$1.04 131149 40.3$\pm$9.1 19.4$\pm$1.24
NCF-$\mathit{t}_{2}$ 308240 1.68$\pm$0.32 1.99$\pm$0.31 131149 3.33$\pm$0.14 2.83$\pm$0.23
SM (×$\times 10^{-3}$) BT (×$\times 10^{-1}$)
@@
CAVIA 50486 979.1$\pm$141.2 859.1$\pm$70.7 116665 21.93$\pm$1.8 22.6$\pm$7.22
CoDA 50547 156.0$\pm$40.52 8.28$\pm$0.29 119679 25.40$\pm$9.5 19.47$\pm$11.6
NCF-$\mathit{t}_{1}$ 50000 680.6$\pm$320.1 677.2$\pm$18.7 117502 21.53$\pm$8.9 20.89$\pm$12.0
NCF-$\mathit{t}_{2}$ 50000 6.42$\pm$0.41 2.03$\pm$0.12 117502 3.46$\pm$0.09 3.77$\pm$0.15
GS (×$\times 10^{-3}$) NS (×$\times 10^{-3}$)
@@
We report in Fig. 3 how well NCF performs on 625 adaptation environments obtained by varying its parameters $\beta$ and $\delta$ on a 25$\times$25 uniform grid. It shows a consistently low MAPE below 2%, except in the bottom right corner where the MAPE rises to roughly 15%, still a remarkably low value for this problem. We highlight the 9 meta-training environments in yellow, and the 4 environments used for OoD adaptation for Table 2 in indigo. Remarkably, our training environment’s convex-hull where adaptation is particularly low is much larger compared to CoDA’s (Kirchmeyer et al., 2022, Figure 3).
@@ 5.3 Conclusion and Future Work
While the benefits of NCFs are evidenced in Section 5.1, its negative impacts should not be neglected. For instance, malicious deployment of such adaptable models in scenarios they were not designed for could lead to serious adverse outcomes. With that in mind, our code, data, and models are openly available at https://github.com/ddrous/ncflow.
@@ A.1 Second-order Taylor expansion with JVPs
Let $x,x_{0}\in\mathbb{R}^{d_{\xi}}$. The second-order Taylor-expansion of $f$ includes its Hessian that we view as a 3-dimensional tensor, and we contract along its last axis such that
@@
[…]
@@ A.2 Identifiability of physical parameters
Identifying the underlying physical parameters of a system is of enduring interest to scientific machine learning practitioners. Our work critically builds on CAMEL Blanke & Lelarge (2024) which extensively studies a model similar to NCF-$\mathit{t}_{1}$ as a linearly parametrized system. Its Proposition 1 states, informally, that in the limit of vanishing training loss $\mathcal{L}(\cdot,\cdot,\cdot)=0$, the relationship between the learned context vectors and the system parameters is linear and can be estimated using ordinary least square. This forces the model to learn a meaningful representation of the system instead of overfitting the examples from the training tasks.
@@
which is valid for all $e\in\operatorname{[\![}1,m\operatorname{]\!]}$. This indicates that the right hand side of Eq. 18 is constant, i.e. $\exists\,\tilde{q}\in\mathbb{R}^{d_{\xi}}$ such that $\forall\,e\in\operatorname{[\![}1,m\operatorname{]\!]}$, $f_{\theta}(\xi^{e})-A\xi^{e}=\tilde{q}$ We then use the fact that in the limit of zero training loss, the predicted and true vector fields coincide for a context $\xi\in U$ and its corresponding underlying parameters $c$ :
@@ A.4 Ordinary alternating minimization
2:$\,\,\theta\in\mathbb{R}^{d_{\theta}}$ randomly initialized
3:$\,\,\xi^{1:m}:=\bigcup\limits_{e=1}^{m}\xi^{e}$, where $\xi^{e}=\mathbf{0}\in\mathbb{R}^{d_{\xi}}$
4:$\,\,\eta_{\theta},\eta_{\xi}>0$
@@
2:$\,\,\theta\in\mathbb{R}^{d_{\theta}}$ learned
3:$\,\,\xi^{a:b}=\bigcup\limits_{e^{\prime}=a}^{b}\xi^{e^{\prime}}$, where $\xi^{e^{\prime}}:=\mathbf{0}\in\mathbb{R}^{d_{\xi}}$
4:$\,\,\eta>0$
@@ A.5 Bulk adaptation algorithm
    Slower convergence: the bulk algorithm could take longer to converge to poorer context vectors when the jointly adapted environments are all very far apart. This is because the contextual self-modulation process would be rendered useless by such task discrepancy, and sampling $j=e$ in the context pool $\mathrm{P}$ to return to standard a Neural ODE in Section 3 would be harder. We find in practice that manually setting $j=e$ by adjusting the vector field to forego the Taylor expansion works well (see Fig. 3). A more direct way of achieving the same result is to disregard $\mathrm{P}$ and retain $f_{\theta}$ (rather than its Taylor expansion) in Eq. 5.
@@ B.1 Gen-Dynamics
$\text{MSE}(x,\hat{x})$ $=\frac{1}{N\times d}\sum_{n=1}^{N}\|x(t_{n})-\hat{x}(t_{n})\|_{2}^{2},$ (21)
$\text{MAPE}(x,\hat{x})$ $=\frac{1}{N\times d}\sum_{n=1}^{N}\left|\frac{x(t_{n})-\hat{x}(t_{n})}{x(t_{n})}\right|\times 100.$ (22)
@@ B.2 Uncertainty Estimation
[…]
@@
where $S$ indicates the number of trajectories used per environment, $N$ the length of each trajectory, and $d$ the dimensionality of the problem. The number of InD environments is denoted by $m$, to be replaced with $b-a+1$ for OoD cases. [^7] Consequently, the first summation symbol’s bounds and its corresponding factor in the denominators of Eqs. 23, 24 and 25 should be adjusted accordingly for OoD formulae. We set $p=m$ for InD uncertainty metrics calculation, and we are guaranteed the existence of those same $m$ environments for OoD cases. However, if our model performs well across all adaptation environments it encounters –as is the case with NCF-$\mathit{t}_{2}$ as observed in Table 2– then all training and adaptation environments can be used, resulting in $p=m+b-a+1$ instead of only $p=m$ environments at our disposal for Eq. 26. This produces more sample predictions, allowing for more reliable population statistics.
@@
To test our hypothesis, we plot the standard deviations as they evolve with time in Fig. 9. When compared to per-problem InD and OoD CLs from Table 3, Fig. 9 reveals that higher confidence levels align with higher standard deviations, which is particularly noticeable on forecasts beyond training time horizons. Additionally, in Fig. 10, we plot the pointwise standard deviations as they relate to the corresponding absolute errors. Naturally, we observe a non-negligible correlation of the two, especially on the GO problem. Focusing on OoD behavior, our model successfully avoids undesirable regions of low uncertainty but higher-than-InD errors (along the lop left corners). Instead, some OoD predictions for ODE problems fall in a region of higher-than-InD uncertainty, but still low error (along the bottom right corners), which stresses the well-suitedness of our approach for OoD generalization. Despite not knowing the underlying source of uncertainty we wish to model, these results suggest that our framework is capable of providing meaningful uncertainty estimates. This said, we emphasize that the calculation and interpretation of aforementioned uncertainty metrics (e.g. the width of CI) should be grounded on knowledge and goals of the problem at hand.
@@ B.3 Simple Pendulum (SP)
$\frac{{\text{d}\alpha}}{{\text{d}t}}&=\omega,\\ \frac{{\text{d}\omega}}{{\text{d}t}}&=-\frac{{\color[rgb]{1,0,0}g}}{L}\sin(\alpha).\end{aligned}\right.$ (28)
@@
We compare our model to the two OFA and OPE baselines as the number of trajectories $S$ in each training environment is increased from 1 to 12. The results reported in Fig. 12 indicate that NCF is indeed the best option when data is limited, while the improvements in OFA MSEs are barely noticeable. As $S$ increses, we observe that OPE is able to overcome overfitting to ultimately achieve the best results. These results demonstrate that NCF effiently uses its few-shots trajectories. However, if neither data nor training time are not a concern (cf. Table 4), then the traditional One-Per-Env should be prioritized.
Figure 12: In-Domain MSEs on the SP problem comparing the sample eficiency of NCF-$\mathit{t}_{1}$ against OPE and OFA. NCF is effective in low-data regimes, and OPE overcomes the gap as data increases.
@@
The computational speed of any method based on Neural ODEs depends on the numerical integrator it uses. To provide consistent number of function evaluations (NFEs), we switch the adaptive time-stepper Dopri5 for the fixed time-stepper RK4, then we measure the duration of epochs as the training progresses. These times as used to produce Fig. 13, which indicates training times per epoch (in seconds) as the number of training environments is increased while keeping the range of gravities unchanged between 2 and 24. We observe excellent scaling, with the training time only increasing by roughly 23% (from 0.38 to 0.47 seconds) when the number of environments is scaled by 10 (from 5 to 50).

### v6

@@ 1 Introduction
    We present an interpretable framework for Multi-Task representation learning, incorporating a straightforward method for uncertainty estimation. This work extends the emerging trend of explainable linearly parameterized physical systems Blanke & Lelarge (2024) to non-linear settings, thus broadening its applicability. For affine systems, we provide a concise proof for the identifiability of their underlying parameters.
@@ 2 Related Work
While LEADS excels at interpolation tasks, it performs poorly during extrapolation Kirchmeyer et al. (2022). Furthermore, it requires retraining a new network $g^{e}_{\psi}$ each time a new environment is encountered, which can be constraining in scenarios where adaptation is frequently required. Before LEADS, other MTL approaches had been proposed outside the context of dynamical systems Caruana (1997); Rebuffi et al. (2017; 2018); Lee & Parish (2021). Still, they do not address the crucial adaptation to new tasks, which is our focus.
@@ 3 Neural Context Flow
We depict the NCF framework in Fig. 2, along with a 3-networks architecture that lifts the contexts and the state vectors into the same representational space before they can interact. Concretely, $x_{i}^{e}(t)$ and the context $\xi$ are first processed independently into $\tilde{x}_{i}^{e}(t)$ (by a state network in blue) and $\tilde{\xi}$ (by a context network in green) respectively, before they are concatenated and fed to a main network (in purple) to produce $f_{\theta}(x_{i}^{e}(t),\xi)$. This explicitly allows the model to account for the potentially nonlinear relationship between the context and the state vector at each evaluation of the vector field.
@@ 3.1 Meta-Training
The above comparison demands the definition of two variants of Neural Context Flows. NCF-$\mathit{t}_{2}$ uses second-order Taylor expansion in equation 5, and is trained via proximal alternating minimization (Algorithm 1). NCF-$\mathit{t}_{1}$ on the other hand, is implemented using a first-order Taylor expansion, and trained using ordinary alternating minimization (Algorithm 3). Although less expressive than $\mathit{t}_{2}, NCF$-$\mathit{t}_{1}$ is faster and serves as a powerful baseline in our experiments.
@@
[…]
@@ 4.1 Experimental Setting
The NCF framework is evaluated on seven seminal benchmarks. The Simple Pendulum (SP) models the periodic motion of a point mass suspended from a fixed point. The Lotka-Volterra (LV) system models the dynamics of an ecosystem in which two species interact. Additional ODEs include a simple model for yeast glycolysis: Glycolytic-Oscillator (GO) Daniels & Nemenman (2015), and the more advanced Sel’kov Model (SM) Strogatz (2018); Sel’Kov (1968). Like GO, SM is non-linearly parameterized, but additionally exhibits starkly different behaviors when its key parameter is varied (see LABEL:fig:attractors). Finally, we consider three PDEs, all with periodic boundary conditions, and cast as ODEs via the method of lines: the non-linear oscillatory Brusselator (BT) model for autocatalytic chemical reactions Prigogine & Lefever (1968), the Gray-Scott (GS) system also for reaction-diffusion in chemical settings Pearson (1993), and Navier-Stokes (NS) for incompressible fluid flow Stokes et al. (1851).
@@ 4.2 Interpolation Results
Train ($\times 10^{-1}$) Adapt ($\times 10^{-3}$)
@@
Table 1 emphasizes the adaptation-time merits of Meta-Learning approaches like NCFs in place of baselines where one context-agnostic vector field is trained for all environments indiscriminately (One-For-All or OFA), or for each environment independently (One-Per-Env or OPE). Additional results and further analysis of these differences is done in Section B.3.
@@ 4.3 Extrapolation Results
A Meta-Learning algorithm is only as good as its ability to extrapolate to unseen environments. In this section we tackle several one-shot generalization problems, i.e. $S^{\prime}=1$. We consider the LV, GO, SM, GS, BT, and NS problems, which all involve adaptation environments outside their meta-training distributions. We compare both variants of NCF to two baselines: CAVIA Zintgraf et al. (2019) which is conceptually the closest GBML method to ours, and CoDA-$\ell_{1}$ Kirchmeyer et al. (2022). Their hyperparameters, laid out in Appendix D, are tuned for a balance of computational efficacy and performance, all the while respecting each baseline’s key assumptions.
At similar parameter counts [^3] This accounts for the total number of learnable parameters in the framework excluding the context vectors, whose size might vary with the method as explained in Appendix D., Table 2 shows that CAVIA is the least effective for learning all six physical systems, with a tendency to overfit on the few shots it receives during both meta-training and meta-testing Mishra et al. (2017). NCF-$\mathit{t}_{2}$ achieves SoTA OoD results on 5 out of 6 problems. While CoDA retains its superiority on GS, we find that it struggles on non-linear problems, most notably on the SM problem whose trajectories go through 3 distinct attractors in a Hopf bifurcation.
@@
[…]
SM ($\times 10^{-3}$) BT ($\times 10^{-1}$)
@@
[…]
GS ($\times 10^{-3}$) NS ($\times 10^{-3}$)
@@
We report in Fig. 3 how well NCF performs on 625 adaptation environments obtained by varying its parameters $\beta$ and $\delta$ on a $\times25$ uniform grid. It shows a consistently low MAPE below 2%, except in the bottom right corner where the MAPE rises to roughly 15%, still a remarkably low value for this problem. We highlight the 9 meta-training environments in yellow, and the 4 environments used for OoD adaptation for Table 2 in indigo. Remarkably, our training environment’s convex-hull where adaptation is particularly low is much larger compared to CoDA’s (Kirchmeyer et al., 2022, Figure 3).
@@
(a) (a) Sample trajectories from the SM problem illustrating the Hopf bifurcation; (b) Losses during meta-training; (c) Subsequent adaptation MSE per method per environment.
@@ 5.3 Conclusion and Future Work
While the benefits of NCFs are evidenced in Section 5.1, their negative impacts should not be neglected. For instance, malicious deployment of such adaptable models in scenarios they were not designed for could lead to serious adverse outcomes. With that in mind, our code, data, and models are openly available at https://github.com/ddrous/ncflow.
@@ A.1 Second-order Taylor expansion with JVPs
Let $x,x_{0}\in\mathbb{R}^{d_{\xi}}$. The second-order Taylor expansion of $f$ includes its Hessian that we view as a 3-dimensional tensor, and we contract along its last axis such that
@@
$=\nabla f(y)(x-y)-\nabla f(y)h+\left[\nabla^{2}f(y)h\right](x-y)+o(\|h\|)+\cancel{o(\|h\|^{2})}$
@@ A.2 Identifiability of physical parameters
Identifying the underlying physical parameters of a system is of enduring interest to scientific machine learning practitioners. Our work critically builds on CAMEL Blanke & Lelarge (2024) which extensively studies a model similar to NCF-$\mathit{t}_{1}$ as a linearly parametrized system. Its Proposition 1 states, informally, that in the limit of vanishing training loss $\mathcal{L}(\cdot,\cdot,\cdot)=0$, the relationship between the learned context vectors and the system parameters is linear and can be estimated using ordinary least squares. This forces the model to learn a meaningful representation of the system instead of overfitting the examples from the training tasks.
@@
which is valid for all $e\in\operatorname{[\![}1,m\operatorname{]\!]}$. This indicates that the right hand side of Eq. 18 is constant, i.e. $\exists\,\tilde{q}\in\mathbb{R}^{d_{\xi}}$ such that $\forall\,e\in\operatorname{[\![}1,m\operatorname{]\!]}$, $f_{\theta}(\xi^{e})-A\xi^{e}=\tilde{q}$ . We then use the fact that in the limit of zero training loss, the predicted and true vector fields coincide for a context $\xi\in U$ and its corresponding underlying parameters $c$ :
@@ A.4 Ordinary alternating minimization
[…]
@@
[…]
@@ A.5 Bulk adaptation algorithm
    Slower convergence: the bulk algorithm could take longer to converge to poorer context vectors when the jointly adapted environments are all very far apart. This is because the contextual self-modulation process would be rendered useless by such task discrepancy, and sampling $j=e$ in the context pool $\mathrm{P}$ to return to a standard Neural ODE in Section 3 would be harder. We find in practice that manually setting $j=e$ by adjusting the vector field to forego the Taylor expansion works well (see Fig. 3). A more direct way of achieving the same result is to disregard $\mathrm{P}$ and retain $f_{\theta}$ (rather than its Taylor expansion) in Eq. 5.
@@ B.1 Gen-Dynamics
[…]
@@ B.2 Uncertainty Estimation
$\text{Rel. MSE}=\frac{100}{m\times S\times N\times d}\sum_{e=1}^{m}\sum_{i=1}^{S}\sum_{n=1}^{N}\frac{\|x^{e}_{i}(t_{n})-\hat{\mu}^{e}_{i}(t_{n})\|_{2}^{2}}{\|x^{e}_{i}(t_{n})\|_{2}^{2}},$ (23)
@@
where $S$ indicates the number of trajectories used per environment, $N$ the length of each trajectory, and $d$ the dimensionality of the problem. The number of InD environments is denoted by $m$, to be replaced with ba+1b-a+1 for OoD cases. [^7] Consequently, the first summation symbol’s bounds and its corresponding factor in the denominators of Eqs. 23, 24 and 25 should be adjusted accordingly for OoD formulae. We set $p=m$ for InD uncertainty metrics calculation, and we are guaranteed the existence of those same $m$ environments for OoD cases. However, if our model performs well across all adaptation environments it encounters –as is the case with NCF-$\mathit{t}_{2}$ as observed in Table 2– then all training and adaptation environments can be used, resulting in p=m+ba+1p=m+b-a+1 instead of only $p=m$ environments at our disposal for Eq. 26. This produces more sample predictions, allowing for more reliable population statistics.
@@
To test our hypothesis, we plot the standard deviations as they evolve with time in Fig. 9. When compared to per-problem InD and OoD CLs from Table 3, Fig. 9 reveals that higher confidence levels align with higher standard deviations, which is particularly noticeable on forecasts beyond training time horizons. Additionally, in Fig. 10, we plot the pointwise standard deviations as they relate to the corresponding absolute errors. Naturally, we observe a non-negligible correlation between the two, especially on the GO problem. Focusing on OoD behavior, our model successfully avoids undesirable regions of low uncertainty but higher-than-InD errors (along the top left corners). Instead, some OoD predictions for ODE problems fall in a region of higher-than-InD uncertainty, but still low error (along the bottom right corners), which stresses the well-suitedness of our approach for OoD generalization. Despite not knowing the underlying source of uncertainty we wish to model, these results suggest that our framework is capable of providing meaningful uncertainty estimates. This said, we emphasize that the calculation and interpretation of aforementioned uncertainty metrics (e.g., the width of CI) should be grounded on knowledge and goals of the problem at hand.
@@ B.3 Simple Pendulum (SP)
$\left\{\begin{aligned} \frac{{\text{d}\alpha}}{{\text{d}t}}&=\omega,\\ \frac{{\text{d}\omega}}{{\text{d}t}}&=-\frac{{\color[rgb]{1,0,0}g}}{L}\sin(\alpha).\end{aligned}\right$. (28)
@@
We compare our model to the two OFA and OPE baselines as the number of trajectories $S$ in each training environment is increased from 1 to 12. The results reported in Fig. 12 indicate that NCF is indeed the best option when data is limited, while the improvements in OFA MSEs are barely noticeable. As $S$ increases, we observe that OPE is able to overcome overfitting to ultimately achieve the best results. These results demonstrate that NCF efficiently uses its few-shot trajectories. However, if neither data nor training time is a concern (cf. Table 4), then the traditional One-Per-Env should be prioritized.
Figure 12: In-Domain MSEs on the SP problem comparing the sample efficiency of NCF-$\mathit{t}_{1}$ against OPE and OFA. NCF is effective in low-data regimes, and OPE overcomes the gap as data increases.
@@
The computational speed of any method based on Neural ODEs depends on the numerical integrator it uses. To provide a consistent number of function evaluations (NFEs), we switch the adaptive time-stepper Dopri5 for the fixed time-stepper RK4, then we measure the duration of epochs as the training progresses. These times as used to produce Fig. 13, which indicates training times per epoch (in seconds) as the number of training environments is increased while keeping the range of gravities unchanged between 2 and 24. We observe excellent scaling, with the training time only increasing by roughly 23% (from 0.38 to 0.47 seconds) when the number of environments is scaled by 10 (from 5 to 50).